OpenStudy (nikkibxox):

How do you solve a system of equations? I don't have a specific equation that needs to be solved I just need help figuring it out. I cant seem to get it. Medal... Fan...Everything... Thanks.

OpenStudy (freckles):

Like x+6=3 ?

OpenStudy (freckles):

oh system?

OpenStudy (freckles):

2 variables?

OpenStudy (nikkibxox):

Like 2x – 3y = –2 4x + y = 24

OpenStudy (freckles):

well there are a couple of ways to go about it: elimination, substitution, ..

OpenStudy (freckles):

Like you can solve the second equation for y. Use that y in the first equation then solve for x.

OpenStudy (freckles):

So first can you isolate y in the following equation: 4x+y=24

OpenStudy (nikkibxox):

y=4x+24?

OpenStudy (freckles):

Remember whatever you do to one side you must do to the other... You subtracted 4x on the left hand side but added 4x on the right hand side.. It should have been you subtracted 4x on both sides.

OpenStudy (freckles):

4x+y=24 subtract 4x on both sides 4x+y-4x=24-4x the whole point of this is to get y on the left hand side by itself 4x-4x=0 so you have y=24-4x also can be written as y=-4x+24

OpenStudy (nikkibxox):

Makes sense.

OpenStudy (freckles):

2x – 3y = –2 now we are going to take that y we just found in terms of x and insert in the first equation 2x-3(-4x+24)=-2 Solve this for x.

OpenStudy (nikkibxox):

One question would i distribute the -3 to the parenthesis? Or just to PEMDAS

OpenStudy (freckles):

you have to distribute the distributive property says a(b+c)=ab+ac then you can apply order of operations with common terms on the left hand side

OpenStudy (nikkibxox):

Got it. 2x-12-72=-2 -10x-72=-2 -10=70 x=-7

OpenStudy (nikkibxox):

Wait. It would be +12

OpenStudy (freckles):

so you mean to say 2x+12x-72=-2 ?

OpenStudy (nikkibxox):

Yes.

OpenStudy (freckles):

ok and this equation 2x+12x-72=-2 is right so far

OpenStudy (nikkibxox):

If I refigure that it would be 2x+12x-72=-2 +72 +72 2x+12x=70 14x=70 x=5

OpenStudy (freckles):

groovy

OpenStudy (freckles):

now remember that one equation that we solved for y in terms of x...?

OpenStudy (freckles):

y=-4x+24 this one!

OpenStudy (nikkibxox):

Yeah. Plug it in?

OpenStudy (freckles):

you can finally find y now by inputting the x you just found into this equation (you could also used the other equation; it just seems easier to plug into this one because we already have y isolated)

OpenStudy (freckles):

and yes so you have y=-4(5)+24

OpenStudy (nikkibxox):

y=4

OpenStudy (freckles):

This method we used was called substitution. Let me show you elimination on this problem.

OpenStudy (freckles):

 This was the system .